OpenStudy (anonymous):

solve 9/11x + 1/x=1 the solution is?? Or is there no solution??

myininaya (myininaya):

clear the fractions by multiplying both sides by the lcm of the bottoms

myininaya (myininaya):

it that first bottom 11x?

OpenStudy (anonymous):

yes :D

myininaya (myininaya):

so multiply both sides by 11x this will get rid of the fractions

OpenStudy (anonymous):

so 9x11 and 1x 11?

myininaya (myininaya):

\[9+11=11x\] 20=11x now to find x divide both sides by 11

OpenStudy (anonymous):

20 doesnt go into 11

OpenStudy (anonymous):

1/11=11

OpenStudy (anonymous):

???

myininaya (myininaya):

20/11=x

OpenStudy (anonymous):

OOOHHH duh!